Two people are in a bicycle race around a circular track. One rider can race completely around the track in 40 seconds. The other rider takes 45 seconds. If they both begin the race at the designated starting point, how long will it take them to be together at this starting point again if they continue to race around the track?
Answer by ankor@dixie-net.com(22740)

Two people are in a bicycle race around a circular track.
One rider can race completely around the track in 40 seconds.
The other rider takes 45 seconds.
If they both begin the race at the designated starting point, how long will it take them to be together at this starting point again if they continue to race around the track?
:
Find the least common multiple of 40 and 45, using prime factoring.
40: 2*2*2*5
45: 3*3*5
LCM 2*2*2*2*3*3*5 = 360 seconds the will both be at the starting point
1st biker: 360/40 = 9 times around the track
2nd biker: 360/45 = 8 times around the track